# CAMA: Enhancing Multimodal In-Context Learning with Context-Aware Modulated Attention

### v2

@@ Conclusion
We introduce CAMA, a plug-and-play and training-free attention modulation method that enhances multimodal ICL. We begin by analyzing the attention dynamics of LVLMs in this setting and, guided by the findings, design a two-stage modulation pipeline for CAMA. CAMA effectively corrects attention deficits and guides the model to focus on image regions that contribute the most to multimodal ICL. Experiments across multiple benchmarks and LVLMs show that CAMA achieves superior results. We believe that CAMA can inspire new directions for future advances of LVLM.
@@ 2   Experiments on Attention Deficits in LVLMs
$I_{i,l}\;=\;\bigl{\lvert}\,\mathbf{A}_{i,l}\;\odot\;\tfrac{\partial\mathcal{L}(s_{i})}{\partial\mathbf{A}_{i,l}}\bigr{\rvert}$, (16)
where $\mathbf{A}_{i,l}$ is the layer’s attention matrix, $\mathcal{L}(s_{i})$ is the loss function for input $s_{i}$, $\odot$ denotes element-wise (Hadamard) product, and the absolute value is applied element-wise. $I(i,j)$ reflects the significance of the information flow from the $j$-th token to the $i$-th token.
@@ 3.1 Setup Details
  • TextVQA: TextVQA contains 34,602 samples in the training set and 5,000 in the validation set. Questions deliberately reference scene text (”What does the billboard say?”), forcing models to integrate OCR output with visual and linguistic cues. The evaluation metric is Accuracy.
  • CLEVR: CLEVR Count Induction (CLEVR), a subset from VL-ICL bench, contains 800 samples in the training set and 200 in the validation set. Each ICD image is paired with an obscure query formatted as an attribute–number pair that identifies specific objects based on four attributes: size, shape, color, or material. We reformat them like ”(Q: Red. A: 2.)”. The task requires models to perform challenging reasoning, first recognizing the task pattern, then counting the correct number of objects. The evaluation metric is Accuracy.
  • MMStar: MMStar consists of 1,500 hand-picked, vision-indispensable samples balanced across 6 core skills and 18 fine-grained axes. Every item was filtered from 22,000 candidates to eliminate data leakage and questions answerable by language alone, so models cannot ”cheat” with memorized knowledge. Since it does not provide a predefined training and validation split, we divide the data using a 7:3 ratio. The evaluation metric is Accuracy, and the reported results are weighted averages across its six subtasks.
  • Flick30k (Young et al. 2014): Flickr30k contains 29,783 samples in the training set and 1,000 in the validation set. It consists of images showing everyday activities, each paired with multiple human-written captions that provide concise descriptions of the scenes. The evaluation metric is CIDEr.
  • MSCOCO (Lin et al. 2014): MSCOCO contains 82,783 samples in the training set and 40,504 in the validation set. It is a widely used benchmark featuring a diverse range of images with detailed and richly descriptive captions, ideal for evaluating a model’s overall perception and description capabilities. The evaluation metric is CIDEr.
  • Hatefulmemes (Kiela et al. 2020): HatefulMemes contains 8,500 samples in the training set and 500 in the validation set. It is designed to reflect real-world challenges found in internet multimodal data. The task requires the model to jointly interpret both the image and the overlaid text to detect instances of hate speech. The evaluation metric is ROC-AUC.
  • L-I-VST (Li et al. 2024b): L-I-VST refers to the Visual StoryTelling subset of LLaVA-Interleave Bench, containing 400 samples. Each sample is an interleaved image-text sequence, where the model is required to generate a new sentence for a given image based on preceding image-sentence pairs, forming a coherent story.
@@
  • +Inst: Instruction augmented (+Inst) prepends to the vanilla sequence a sentence that explicitly instructs the model to carry out ICL: ”First, study the examples we provide. Then utilize what you have learned to answer the new question.” Notably, this sentence achieves the best overall performance among the three styles we tested and is therefore adopted in our main experiments. The other two variants are ”Please make full use of the provided examples to solve a new image-related question” and ”Here are several example image-text pairs, along with a new image and question. Please perform in-context learning.”
  • CD: Contrastive decoding (CD) constructs a distorted variant targeted to a selected part of the input, separately feeds the original and distorted sequences into the model, and collects their final output logits. By calibrating the vanilla logits with those from the distorted input, we obtain new logits whose distribution implicitly ensures that the model attends to the selected part. In LVLMs, CD is typically employed to encourage the model to rely more on visual evidence and thus mitigate hallucination or language prior. A common practice is to either add noise to the input image or directly replace it with a blank image. CD can likewise enhance multimodal ICL. We observe that replacing the images with blank images yields a larger improvement than adding noise. After calibration, the resulting logits accentuate the influence of the ICDs. Formally, CD is presented by:
@@
    where $\theta$ denotes the LVLM, $y_{t}$ denotes the model output at time step $t$, and $X$ and $\hat{X}$ denote the vanilla sequence and the distorted sequence, respectively. $\alpha$ is a user-defined value. For all VQA benchmarks, we set $\alpha=0.4$. For other benchmarks, we set $\alpha=0.6$.
@@ 3.4 Additional Qualitative Visualizations
In Figure 8, we present the ICL enhancement results of CAMA for four input sequences by visualizing the image attention in the 19th decoder layer of LLaVA-NeXT-7B. In (a), the sequence comes from CLEVR. This benchmark poses a challenge for the model to identify the task mapping in ICL because of its subtle textual descriptions. Through dynamic attention increments, CAMA highlights the key visual features that align with the text semantics, helping the model infer the correct task mapping. As shown in the figure, the model accurately locates the objects whose attributes match those in the question and successfully performs the counting task. In (b), the example illustrates a typical VQA scenario. In this sequence, although the third ICD and the query sample both depict cakes as the main subject, the first ICD is actually more crucial for solving the query sample, as both focus on identifying food ingredients. CAMA effectively prevents the first ICD from being overlooked due to position bias, guiding the model’s attention toward the connection between ”drink” and ”strawberry” in the first ICD. This association helps the model identify the relationship between ”cake” and ”banana” in the query sample, enabling it to provide the correct answer. In (c), the example is from L-I-VST, which requires the model to complete an interleaved story. While reinforcing the alignment between each image–text pair, CAMA also guides the model to focus on consistent visual cues between earlier images and the query sample image, such as ”a girl wearing pink clothes,” thereby recognizing the narrative as a coherent story. This enables the model to generate text that follows the format established in the preceding context. In (d), the sequence comes from MMStar, which presents tasks in a multiple-choice format and requires the LVLM to recognize the emotion in an image. CAMA identifies the second ICD as the most helpful for answering the query sample, since both images are cat memes. The perception of the second image is thus enhanced, allowing it to fully capture the connection between the cat’s expression and its ground-truth option ”sadness.” When interpreting the query sample image, the model focuses more on the cat’s expression rather than the text in the image, demonstrating CAMA’s ability to deeply promote effective multimodal ICL. However, in our experiments we also find that CAMA struggles when the in-context sequence contains too much irrelevant information, such as cases where seven out of eight or all ICDs are unrelated to the query sample or provide misleading content. In these situations, the difficulty of appropriately allocating attention in Stage II prevents the model’s behavior from being steered in the desired direction. These failure cases highlight that improving the quality of sequence configuration remains important.

### v3

@@ Conclusion
We introduce CAMA, a training-free and plug-and-play attention modulation method that enhances multimodal ICL. We begin by analyzing the attention dynamics of LVLMs in this setting and, guided by the findings, design a two-stage modulation pipeline for CAMA. CAMA effectively corrects attention deficits and guides the model to focus on image regions that contribute the most to multimodal ICL. Experiments across multiple benchmarks and LVLMs show that CAMA achieves superior results. We believe that CAMA can inspire new directions for future advances of LVLM.
@@
Acknowledgments
We acknowledge the computing resources provided by NSF ACCESS.
@@ 2   Experiments on Attention Deficits in LVLMs
$I_{i,l}\;=\;\bigl\lvert\,A_{i,l}\;\odot\;\tfrac{\partial\mathcal{L}(s_{i})}{\partial\mathbf{A}_{i,l}}\bigr\rvert$, (16)
where $A_{i,l}$ is the post-softmax attention matrix of the layer, $\mathcal{L}(s_{i})$ is the loss function for input $s_{i}$, $\odot$ denotes element-wise (Hadamard), and the absolute value is applied element-wise. $I(i,j)$ reflects the significance of the information flow from the $j$-th token to the $i$-th token.
@@ 3.1 Setup Details
  • TextVQA: TextVQA contains 34,602 samples in the training set and 5,000 in the validation set. Questions deliberately reference scene text (“What does the billboard say?”), forcing models to integrate OCR output with visual and linguistic cues. The evaluation metric is Accuracy.
  • CLEVR: CLEVR Count Induction (CLEVR), a subset from VL-ICL bench, contains 800 samples in the training set and 200 in the validation set. Each ICD image is paired with an obscure query formatted as an attribute–number pair that identifies specific objects based on four attributes: size, shape, color, or material. We reformat them like “(Q: Red. A: 2.)”. The task requires models to perform challenging reasoning, first recognizing the task pattern, then counting the correct number of objects. The evaluation metric is Accuracy.
  • MMStar: MMStar consists of 1,500 hand-picked, vision-indispensable samples balanced across 6 core skills and 18 fine-grained axes. Every item was filtered from 22,000 candidates to eliminate data leakage and questions answerable by language alone, so models cannot “cheat” with memorized knowledge. Since it does not provide a predefined training and validation split, we divide the data using a 7:3 ratio. The evaluation metric is Accuracy, and the reported results are weighted averages across its six subtasks.
  • Flickr30k (flickr): Flickr30k contains 29,783 samples in the training set and 1,000 in the validation set. It consists of images showing everyday activities, each paired with multiple human-written captions that provide concise descriptions of the scenes. The evaluation metric is CIDEr.
  • MSCOCO (mscoco): MSCOCO contains 82,783 samples in the training set and 40,504 in the validation set. It is a widely used benchmark featuring a diverse range of images with detailed and richly descriptive captions, ideal for evaluating a model’s overall perception and description capabilities. The evaluation metric is CIDEr.
  • Hatefulmemes (hateful): HatefulMemes contains 8,500 samples in the training set and 500 in the validation set. It is designed to reflect real-world challenges found in internet multimodal data. The task requires the model to jointly interpret both the image and the overlaid text to detect instances of hate speech. The evaluation metric is ROC-AUC.
  • L-I-VST (llava): L-I-VST refers to the Visual StoryTelling subset of LLaVA-Interleave Bench, containing 400 samples. Each sample is an interleaved image-text sequence, where the model is required to generate a new sentence for a given image based on preceding image-sentence pairs, forming a coherent story.
@@
  • +Inst: Instruction augmented (+Inst) prepends to the vanilla sequence a sentence that explicitly instructs the model to carry out ICL: “First, study the examples we provide. Then utilize what you have learned to answer the new question.” Notably, this sentence achieves the best overall performance among the three styles we tested and is therefore adopted in our main experiments. The other two variants are “Please make full use of the provided examples to solve a new image-related question” and “Here are several example image-text pairs, along with a new image and question. Please perform in-context learning.” (h5)
  • CD: Contrastive decoding (CD) constructs a distorted variant targeted to a selected part of the input, separately feeds the original and distorted sequences into the model, and collects their final output logits. By calibrating the vanilla logits with those from the distorted input, we obtain new logits whose distribution implicitly ensures that the model attends to the selected part. In LVLMs, CD is typically employed to encourage the model to rely more on visual evidence and thus mitigate hallucination or language prior (h1; h2). A common practice is to either add noise to the input image or directly replace it with a blank image. CD can likewise enhance multimodal ICL. We observe that replacing the images with blank images yields a larger improvement than adding noise. After calibration, the resulting logits accentuate the influence of the ICDs. Formally, CD is presented by:
@@
    where $\theta$ denotes the LVLM, $y_{t}$ denotes the model output at time step $t$, and $X$ and $\hat{X}$ denote the vanilla sequence and the distorted sequence, respectively. $\alpha$ is a user-defined value. For all VQA benchmarks, we set $\alpha=0.4$. For other benchmarks, we set $\alpha=0.6$ (h3; h4).
@@ 3.4 Additional Qualitative Visualizations
In Figure 8, we present the ICL enhancement results of CAMA for four input sequences by visualizing the image attention in the 19th decoder layer of LLaVA-NeXT-7B. In (a), the sequence comes from CLEVR. This benchmark poses a challenge for the model to identify the task mapping in ICL because of its subtle textual descriptions. Through dynamic attention increments, CAMA highlights the key visual features that align with the text semantics, helping the model infer the correct task mapping. As shown in the figure, the model accurately locates the objects whose attributes match those in the question and successfully performs the counting task. In (b), the example illustrates a typical VQA scenario. In this sequence, although the third ICD and the query sample both depict cakes as the main subject, the first ICD is actually more crucial for solving the query sample, as both focus on identifying food ingredients. CAMA effectively prevents the first ICD from being overlooked due to position bias, guiding the model’s attention toward the connection between “drink” and “strawberry” in the first ICD. This association helps the model identify the relationship between “cake” and “banana” in the query sample, enabling it to provide the correct answer. In (c), the example is from L-I-VST, which requires the model to complete an interleaved story. While reinforcing the alignment between each image–text pair, CAMA also guides the model to focus on consistent visual cues between earlier images and the query sample image, such as “a girl wearing pink clothes,” thereby recognizing the narrative as a coherent story. This enables the model to generate text that follows the format established in the preceding context. In (d), the sequence comes from MMStar, which presents tasks in a multiple-choice format and requires the LVLM to recognize the emotion in an image. CAMA identifies the second ICD as the most helpful for answering the query sample, since both images are cat memes. The perception of the second image is thus enhanced, allowing it to fully capture the connection between the cat’s expression and its ground-truth option “sadness.” When interpreting the query sample image, the model focuses more on the cat’s expression rather than the text in the image, demonstrating CAMA’s ability to deeply promote effective multimodal ICL. However, in our experiments we also find that CAMA struggles when the in-context sequence contains too much irrelevant information, such as cases where seven out of eight or all ICDs are unrelated to the query sample or provide misleading content. In these situations, the difficulty of appropriately allocating attention in Stage II prevents the model’s behavior from being steered in the desired direction. These failure cases highlight that improving the quality of sequence configuration remains important.